# Boosting Scientific Error-Bounded Lossy Compression through Optimized Synergistic Lossy-Lossless Orchestration

### v1

@@ 5.1. Improving interpolation-based lossy data prediction
Because lossless data encoding algorithms mainly leverage 2 types of data information: token-wise entropy and bit-wise redundancy, there are also 2 direct solutions for improving the compressibility of quantization codes from the cuSZ-Hi interpolation-based data predictor. First, we present better data prediction accuracy and anchor-point rate, making the distribution of quantization codes more concentrated to zero, delivering lower entropy. Second, based on the characteristics of interpolation-based data predictors, we reorder the quantization code array with a fixed mapping, bringing similar values together to raise its compressibility.
@@ 5.1.1. Reorganized data and anchor partitions
In Figure 3, we present the details of cuSZ-Hi thread-block-wise data partition, anchor layout, and interpolation workloads in a 3D example, comparing it to cuSZ-I. Specifically: 1) In cuSZ-Hi, the 3D partitioned data block to be processed by each thread block has a size of $17^{3}$ including all anchors) instead of $33\times 9\times 9$ for cuSZ-I. 2) On each dimension, the stride of losslessly stored anchor points is 16 in cuSZ-Hi instead of $8$ in cuSZ-I. Due to this, the interpolations for 3D data blocks in cuSZ-Hi follow a 4-level hierarchy instead of the cuSZ-I 3-level interpolation. We justify our new design as follows: First, a dimensionally isotropic data partition can be better adaptive to different interpolation computation orders, which will be detailed in Section 5. Second, the data block size in cuSZ-Hi is double the original data block size in cuSZ-I, which better exploits the shared memory and cache in GPU architectures and brings more available data points in the prediction process for improving the data prediction accuracy (Liu et al., 2024a). Moreover, in our preliminary experiments, further enlarging the data block will bring limited compression ratio improvement while significantly reducing the compression speed. Last, according to preliminary experiments, anchor points in cuSZ-I occupy a considerable portion of the compression data size. Reducing the number of anchor points (to $\frac{1}{8}$ of cuSZ-I) can effectively reduce the anchor storage overhead meanwhile preserving the data prediction quality.
@@ 5.1.2. Parallelized multi-dimensional interpolation
Regarding the GPU-parallelized design of interpolation-based data prediction, cuSZ-Hi also introduces a new scheme, which is featured in Figure 4. To further leverage the data correlation along multiple data dimensions, adopting the design concepts proposed in (Liu et al., 2024a), cuSZ-Hi crafts a new parallelized, multi-dimensional-interpolation-based prediction scheme in each partitioned data block (Figure 4 (b)). Unlike original 1D interpolation dimension sequences ((Figure 4 (a)), on each level, parallelized multi-dimensional interpolations follow a process with an increasing number of dimensions used in the interpolation (1D $\to$ 2D $\to$ 3D). This scheme is isotropic along all dimensions (so free of dimension sequence selection) and better leverages the multi-dimensional data correlation than the existing one-dimensional spline interpolation scheme. In predicting one certain data point along multiple dimensions, the prediction value is the interpolation-prediction average along those dimensions. Moreover, to improve the prediction quality, only prediction values with the highest spline order will be used and averaged in the prediction. For example, two cubic spline interpolations along two dimensions in cuSZ-Hi will be averaged as the final data prediction, but if one is cubic and the other is linear/quadratic, cuSZ-Hi will only use the cubic spline interpolation as the prediction result.
@@ 5.1.3. Workload-balanced auto-tuning
cuSZ-Hi introduced multiple interpolation splines and schemes, which adapt differently to input data characteristics. To this end, an online dynamic selection of those splines and schemes will be critical in optimizing the compression ratio of cuSZ-Hi on diverse data inputs. Previously, cuSZ-I merely conducted fast profiling on sampled input data points to determine the best-fit interpolation spline and scheme, which is insufficient to make an accurate selection for interpolation configuration. Following the design proposed in (Liu et al., 2022; Liu et al., 2024a), the compression pipeline of cuSZ-Hi includes a full-functional module of interpolation auto-tuning, which is still lightweight enough, with well-balanced workloads on thread blocks to minimize the overhead in GPU-parallelized compression computing. Specifically, cuSZ-Hi first uniformly samples data blocks (with the same size as per-thread-block-data), with a total volume of $0.2\%$ of the whole data. In the auto-tuning kernel, each sampled data block is distributed to multiple thread blocks to perform interpolations on different levels with different splines and schemes.
To optimize the performance of the auto-tuning kernel, best balancing the number of thread blocks used and per-block workload, cuSZ-Hi manages the interpolation test differently on different levels. For level 4 and level 3 interpolations, which have limited computational cost, interpolation tests with all possible configurations (spline $+$ scheme) are performed on the same threadblock. For level 2 interpolation, the tests are equally allocated to 2 thread blocks. For level 1 interpolation, which requires intensive computation, cuSZ-Hi leverages 6 thread blocks to test one configuration on each. After performing the interpolation tests on each level and with each configuration, cuSZ-Hi collects the aggregated data prediction errors from each test and selects the interpolation configuration separately on each level that has the minimized test error.
@@ 5.1.4. Mapping-based quantization codes reordering
The Huffman encoding module (which is adopted by SZ3, cuSZ, and cuSZ-I) is entropy-based, which means that its compression ratios on lossy-predicted quantization codes are solely dependent on the frequency distribution of the tokens (i.e., values of the codes). However, it cannot exploit the sequential order of the tokens and the bit-wise redundancies in the tokens. To this end, cuSZ-Hi incorporates additional bitwise-redundancy-based lossless encoding modules (detailed in Section 5.2), and the quantization codes are also reversibly reordered to optimize the de-redundancy encoding.
@@
As previously mentioned, in cuSZ-Hi, the data predictions are performed by spline interpolations with diverse strides. For example, data points whose coordinates have at least one odd value are predicted by interpolation with stride 1, and the rest of the data are predicted by interpolations with larger strides (2 to 8). In existing cuSZ-I, the quantized prediction errors are gathered in an array that has the same layout as the input data, and the coordinates of each element remain the same as the corresponding input data point. Therefore, when the quantization code array is flattened to a 1D sequence for lossless encoding, the quantization codes from different interpolation levels and strides will get shuffled in the sequence. However, several existing research works (Liu et al., 2022; Liu et al., 2024a, c) have observed and proved that the prediction accuracy of interpolation-based data predictor is highly relevant to the interpolation stride, for which larger interpolation strides lead to higher data prediction errors. Therefore, quantization codes from large-stride interpolations often present larger absolute values and exhibit different distributions from quantization codes generated by small-stride interpolations. Those insights into the distribution of the quantization codes drive us to rearrange the to-be-encoded quantization code sequence by separating quantization codes from different interpolation strides.
@@
Specifically, after acquiring the multi-dimension quantization code array (the coordinates are the same as the data array), cuSZ-Hi maps the array to a reordered 1D sequence for lossless encoding, in which quantization codes for each interpolation level are grouped in adjacent indices. For example, for a quantization code with coordinate $(x,y,z)$ in the 3D code array, its index $I(x,y,z)$ in the mapped 1D sequence will be ($A$ is the stride of anchor points, like 16, $(d_{x},d_{y},d_{z})$ are the global data dimension sizes):
(3) $\begin{split}I(x,y,z)=&\operatorname{Prefix}(l)+xd_{y}d_{z}+yd_{z}+z\\ -\left\lfloor\frac{x+1}{2}\right\rfloor\left\lfloor\frac{d_{y}+1}{2}\right\rfloor\left\lfloor\frac{d_{z}+1}{2}\right\rfloor&-\mathbf{1}_{\{x\,\text{even}\}}\left\lfloor\frac{y+1}{2}\right\rfloor\left\lfloor\frac{d_{z}+1}{2}\right\rfloor-\mathbf{1}_{\{x,y\,\text{even}\}}\left\lfloor\frac{z+1}{2}\right\rfloor\\ l=\max&\left\{l\in\mathbb{N},0\leq l\leq\log_{2}{A},2^{l}|x,y,z\right\},\\ \operatorname{Prefix}(l)=f_{x}(l)&f_{y}(l)f_{z}(l)\ (l<\log_{2}{A}),\ \operatorname{Prefix}(l)=0\ (l=\log_{2}{A}),\\ f_{x}(0)=\left\lfloor\frac{d_{x}+1}{2}\right\rfloor,&\ f_{y}(0)=\left\lfloor\frac{d_{y}+1}{2}\right\rfloor,\ f_{z}(0)=\left\lfloor\frac{d_{z}+1}{2}\right\rfloor,\\ f_{x}(k+1)=\left\lfloor\frac{f_{x}(k)+1}{2}\right\rfloor,&\ f_{y}(k+1)=\left\lfloor\frac{f_{y}(k)+1}{2}\right\rfloor,\ f_{z}(k+1)=\left\lfloor\frac{f_{z}(k)+1}{2}\right\rfloor\end{split}$
@@
In Eq. 3, $l$ indicates the interpolation level for predicting the data point with coordinate $(x,y,z)$, and $\operatorname{Prefix}(l)$ is the number of data points on higher interpolation levels. According to Eq. 3, quantization codes from the larger interpolation strides will appear first in the 1D code sequence (with smaller sequence index $I(x,y,z)$), followed by codes from smaller interpolation strides. In Figure 5, we present the distribution of quantization code values by indices in the code sequences, both from direct flattening and cuSZ-Hi reordering. In the non-reordered code sequence, the code values oscillate intensively over a wide range. Quantization code reordering transfers most outlier values to a short range (beginning of the sequence), highly increasing the overall smoothness of the sequence to improve the compression ratio.
@@ 5.2. Optimizing lossless ratio and throughput
Past research over years (Liu et al., 2021b; Tian et al., 2021) and recently proposed (Chen et al., 2024; Azami et al., 2025) have proved that the evolving lossless data encoding techniques have great potential in improving the scientific lossy compression over various aspects, including both compression ratio and throughput. According to existing work (Liu et al., 2024c) and our investigation, one fact lies in our findings: Multiple existing scientific lossy compressors have not fully utilized the correlation information in the original data, leaving undealt compressibilities in the output data. In Table 1, we present a showcase, listing the lossless compression ratio by NVIDIA Bitcomp (nvc, [n.d.]) of several compressed outputs from scientific error-bounded lossy compressors, including baselines (detailed in Section 2 and Section 6.1) and our proposed cuSZ-Hi. Most of the compression outputs can still be highly compressed by NVIDIA Bitcomp, to $10\%\sim 40\%$ of the original sizes, and this is just the reason why cuSZ-I directly integrated NVIDIA Bitcomp as a module of its pipeline. Unfortunately, Bitcomp is NVIDIA-proprietary, so integrating it into the compression framework will bring severe restrictions on its extension, customization, and (cross-platform) deployment. To maximize the scientific lossy compression ratio on GPUs, achieving minimal information redundancy while preserving the usability of the whole compression framework, as cuSZ-Hi does, we need a new solution for the lossless pipeline, which will be detailed step-by-step in the next.
@@ 5.2.1. Minimizing quantization code length
The first important design detail of cuSZ-Hi lossless pipeline is that, in the integer quantization code sequence, each value has been resized to the minimized one-byte-width, processed in the uint8_t format (the out-of-range outliers will be collected and stored separately). Different from other scientific lossy compressors that quantize data offsets (Huang et al., 2023, 2024a) or inaccurate extrapolation errors (Tian et al., 2020), interpolation-based data compressors usually produce quantization codes with a highly-concentrated distribution, including very few large values (Liu et al., 2024c). Consequently, applying a short data format to store the quantization codes will accelerate the encoding speed and simplify bit patterns without high overheads for recording outliers.
@@ 5.2.2. Investigation and benchmarking of lossless encoders
In our benchmarking, the evaluated lossless pipelines are collected as follows: First, incorporating Huffman encoding as a ”preprocessor” in the lossless compression pipeline has been a widely adopted scheme in existing compressors (Liang et al., 2023; Liu et al., 2024a, c), so in addition to the standalone lossless modules, we also include a set of Huffman-incorporated variants corresponding to each of them. Second, for the LC framework, we perform preliminary experiments on several datasets and select 8 representative and adaptive pipelines (RRE1, RRE1-RRE2, etc. as shown in Figure 6) with 1/2/3/4 stages (as indicated by Figure 6, pipelines with more stages are not necessary). The benchmarking is conducted on 4 scientific datasets (Zhao et al., 2020a) on a computing platform facilitated with NVIDIA RTX 6000 Ada GPUs.
The compression ratio and overall (compression-decompression) throughput of each lossless pipeline are shown in Figure 6, with Pareto frontiers marked (solutions with throughput less than 25 GiB/s are not included in the frontiers because they bring unacceptable time overheads to the full compression pipeline). Regarding open-source solutions with acceptable compression throughput, we find that the combination of Huffman encoding and the RRE4-TCMS8-RZE1 pipeline achieves high compression ratios on cuSZ-Hi quantization codes. The Zstd (Collet, 2015) compressor in NVIDIA nvCOMP achieves the highest compression ratio. However, not only is this implementation NVIDIA-proprietary, but it also consists of multiple self-contained encoders with coarse component configuration granularity. Consequently, its throughput is significantly low in most scenarios, making it impractical for cuSZ-Hi. The nvCOMP::ANS/Bitcomp and GPULZ may present decent compression ratios in some scenarios, but they show no advantages over the multi-stage LC pipelines. The GDeflate/LZ4/ndzip/Huffman-only lossless solutions exhibit very poor compression ratios and/or throughputs, being infeasible for cuSZ-Hi usage.
According to the benchmarking results and analysis, we adopt the HF-RRE4-TCMS8-RZE1 lossless pipeline in cuSZ-Hi, forming its compression-ratio-preferred mode, cuSZ-Hi-CR. However, Huffman encoding on GPUs also brings a performance bottleneck in the compression framework and lowers the overall compression throughput (Rivera et al., 2022). To this end, regarding Figure 6, cuSZ-Hi integrates an alternative Huffman-encoding-free lossless pipeline, which is the TCMS1-BIT1-RRE1 pipeline with high throughput and decent compression ratio. In Section 6, we will see that, cuSZ-Hi with this lossless pipeline (the throughput-preferred cuSZ-Hi-TP) also gains very acceptable compression ratios with significantly improved compression throughputs, proving its practical usability.

### v2

@@ 5.1. Improving interpolation-based lossy data prediction
Because lossless data encoding algorithms mainly leverage 2 types of data information: token-wise entropy and bit-wise redundancy, there are also 2 direct solutions for improving the compressibility of quantization codes from the cuSZ-Hi interpolation-based data predictor. First, we present improved data prediction accuracy and anchor-point rate, resulting in a more concentrated distribution of quantization codes around zero, which delivers lower entropy. Second, based on the characteristics of interpolation-based data predictors, we reorder the quantization code array using a fixed mapping, grouping similar values to enhance their compressibility.
@@ 5.1.1. Reorganized data and anchor partitions
In Figure 3, we present the details of cuSZ-Hi thread-block-wise data partition, anchor layout, and interpolation workloads in a 3D example, comparing it to cuSZ-I. Specifically: 1) In cuSZ-Hi, the 3D partitioned data block to be processed by each thread block has a size of $17^{3}$, including all anchors) instead of $33\times 9\times 9$ for cuSZ-I. 2) On each dimension, the stride of losslessly stored anchor points is 16 in cuSZ-Hi instead of $8$ in cuSZ-I. Due to this, the interpolations for 3D data blocks in cuSZ-Hi follow a 4-level hierarchy instead of the cuSZ-I 3-level interpolation. We justify our new design as follows: First, a dimensionally isotropic data partition can be better adaptive to different interpolation computation orders, which will be detailed in §5. Second, the data block size in cuSZ-Hi is double the original data block size in cuSZ-I, which better exploits the shared memory and cache in GPU architectures and brings more available data points in the prediction process for improving the data prediction accuracy (Liu et al., 2024a). Moreover, in our preliminary experiments, further enlarging the data block yields limited improvement in compression ratio, while significantly reducing the compression speed. Last, according to preliminary experiments, anchor points in cuSZ-I occupy a considerable portion of the compression data size. Reducing the number of anchor points (to $\frac{1}{8}$ of cuSZ-I) can effectively reduce the anchor storage overhead while preserving data prediction quality.
@@ 5.1.2. Parallelized multi-dimensional interpolation
Regarding the GPU-parallelized design of interpolation-based data prediction, cuSZ-Hi also introduces a new scheme, which is featured in Figure 4. To further leverage the data correlation along multiple data dimensions, adopting the design concepts proposed in (Liu et al., 2024a), cuSZ-Hi crafts a new parallelized, multi-dimensional-interpolation-based prediction scheme in each partitioned data block (Figure 4 (b)). Unlike original 1D interpolation dimension sequences ((Figure 4 (a)), on each level, parallelized multi-dimensional interpolations follow a process with an increasing number of dimensions used in the interpolation (1D $\to$ 2D $\to$ 3D). This scheme is isotropic along all dimensions (free of dimension sequence selection) and better leverages the multidimensional data correlation than the existing one-dimensional spline interpolation scheme. When predicting a specific data point along multiple dimensions, the prediction value is the average of the interpolation predictions along those dimensions. Moreover, to improve the prediction quality, only prediction values with the highest spline order will be used and averaged in the prediction. For example, two cubic spline interpolations along two dimensions in cuSZ-Hi will be averaged as the final data prediction, but if one is cubic and the other is linear/quadratic, cuSZ-Hi will only use the cubic spline interpolation as the prediction result.
@@ 5.1.3. Workload-balanced auto-tuning
cuSZ-Hi introduces multiple interpolation splines and schemes, which adapt differently to the characteristics of the input data. To this end, an online dynamic selection of those splines and schemes will be critical in optimizing the compression ratio of cuSZ-Hi on diverse data inputs. Previously, cuSZ-I merely conducted fast profiling on sampled input data points to determine the best-fit interpolation spline and scheme, which is insufficient to make an accurate selection for interpolation configuration. Following the design proposed in (Liu et al., 2022; Liu et al., 2024a), the compression pipeline of cuSZ-Hi includes a full-functional module of interpolation auto-tuning, which is still lightweight enough, with well-balanced workloads on thread blocks to minimize the overhead in GPU-parallelized compression computing. Specifically, cuSZ-Hi first uniformly samples data blocks (with the same size as per-thread-block-data), with a total volume of $0.2\%$ of the whole data. In the auto-tuning kernel, each sampled data block is distributed to multiple thread blocks to perform interpolations at different levels using various splines and schemes.
To optimize the performance of the auto-tuning kernel, balancing the number of thread blocks used and the per-block workload, cuSZ-Hi manages the interpolation test differently at different levels. For level 4 and level 3 interpolations, which have limited computational cost, interpolation tests with all possible configurations (spline ++ scheme) are performed on the same threadblock. For level 2 interpolation, the tests are equally allocated to 2 thread blocks. For level 1 interpolation, which requires intensive computation, cuSZ-Hi leverages 6 thread blocks to test one configuration on each. After performing the interpolation tests on each level and with each configuration, cuSZ-Hi collects the aggregated data prediction errors from each test and selects the interpolation configuration separately on each level that has the minimized test error.
@@ 5.1.4. Mapping-based quantization codes reordering
The Huffman encoding module (which is adopted by SZ3, cuSZ, and cuSZ-I) is entropy-based, which means that its compression ratios on lossy-predicted quantization codes are solely dependent on the frequency distribution of the tokens (i.e., values of the codes). However, it cannot exploit the sequential order of the tokens and the bit-wise redundancies in the tokens. To this end, cuSZ-Hi incorporates additional bitwise-redundancy-based lossless encoding modules (detailed in §5.2), and the quantization codes are also reversibly reordered to optimize the de-redundancy encoding.
@@
As previously mentioned, in cuSZ-Hi, the data predictions are performed by spline interpolations with diverse strides. For example, data points whose coordinates have at least one odd value are predicted by interpolation with a stride of 1, and the rest of the data are predicted by interpolation with larger strides (2 to 8). In existing cuSZ-I, the quantized prediction errors are gathered in an array that has the same layout as the input data, and the coordinates of each element remain the same as the corresponding input data point. Therefore, when the quantization code array is flattened to a 1D sequence for lossless encoding, the quantization codes from different interpolation levels and strides will get shuffled in the sequence. However, several existing research works (Liu et al., 2022; Liu et al., 2024a, c) have observed and proved that the prediction accuracy of an interpolation-based data predictor is highly relevant to the interpolation stride, for which larger interpolation strides lead to higher data prediction errors. Therefore, quantization codes from large-stride interpolations often present larger absolute values and exhibit different distributions from quantization codes generated by small-stride interpolations. Those insights into the distribution of the quantization codes drive us to rearrange the to-be-encoded quantization code sequence by separating quantization codes from different interpolation strides.
@@
Specifically, after acquiring the multi-dimensional quantization code array (the coordinates are the same as the data array), cuSZ-Hi maps the array to a reordered 1D sequence for lossless encoding, in which quantization codes for each interpolation level are grouped in adjacent indices. $A$ is the stride of anchor points, like 16, $(d_{x},d_{y},d_{z})$ are the global data dimension sizes. For a quantization code with coordinate $(x,y,z)$ in the 3D code array, its index $I(x,y,z)$ in the mapped 1D sequence will be:
[…]
with
(4) $f_{x}(0)=\left\lfloor\frac{d_{x}+1}{2}\right\rfloor,$ $\ f_{x}(k+1)=\left\lfloor\frac{f_{x}(k)+1}{2}\right\rfloor,$
$f_{y}(0)=\left\lfloor\frac{d_{y}+1}{2}\right\rfloor,$ $\ f_{y}(k+1)=\left\lfloor\frac{f_{y}(k)+1}{2}\right\rfloor,$
$f_{z}(0)=\left\lfloor\frac{d_{z}+1}{2}\right\rfloor,$ $\ f_{z}(k+1)=\left\lfloor\frac{f_{z}(k)+1}{2}\right\rfloor.$
@@
In Eq. 3, $\ell$ indicates the interpolation level for predicting the data point with coordinate $(x,y,z)$, and $\operatorname{prefix}(\ell)$ is the number of data points on higher interpolation levels. According to Eq. 3, quantization codes from the larger interpolation strides will appear first in the 1D code sequence (with smaller sequence index $I(x,y,z)$), followed by codes from smaller interpolation strides. In Figure 5, we present the distribution of quantization code values by indices in the code sequences, both from direct flattening and cuSZ-Hi reordering. In the non-reordered code sequence, the code values oscillate intensively over a wide range. Quantization code reordering transfers most outlier values to a short range (at the beginning of the sequence), thereby significantly increasing the overall smoothness of the sequence and improving the compression ratio.
@@ 5.2. Optimizing lossless ratio and throughput
Past research over years (Liu et al., 2021b; Tian et al., 2021) and recently proposed (Chen et al., 2024; Azami et al., 2025) have proved that the evolving lossless data encoding techniques have great potential in improving the scientific lossy compression over various aspects, including both compression ratio and throughput. According to existing work (Liu et al., 2024c) and our investigation, one key finding is that Multiple existing scientific lossy compressors have not fully utilized the correlation information in the original data, leaving undealt compressibilities in the output data. In Table 1, we present a showcase, listing the lossless compression ratio by NVIDIA Bitcomp (nvc, [n.d.]) of several compressed outputs from scientific error-bounded lossy compressors, including baselines (detailed in §2 and §6.1) and our proposed cuSZ-Hi. Most of the compression outputs can still be highly compressed by NVIDIA Bitcomp, reducing their sizes to $10\%\sim 40\%$ of the original size, which is the reason why cuSZ-I directly integrated NVIDIA Bitcomp as a module of its pipeline. Unfortunately, Bitcomp is NVIDIA-proprietary, so integrating it into the compression framework will bring severe restrictions on its extension, customization, and (cross-platform) deployment. To maximize the scientific lossy compression ratio on GPUs, achieving minimal information redundancy while preserving the usability of the whole compression framework, as cuSZ-Hi does, we need a new solution for the lossless pipeline, which will be detailed step-by-step in the next.
@@ 5.2.1. Minimizing quantization code length
The first important design detail of cuSZ-Hi lossless pipeline is that, in the integer quantization code sequence, each value has been resized to the minimized one-byte-width, processed in the uint8_t format (the out-of-range outliers will be collected and stored separately). Unlike other scientific lossy compressors that quantize data offsets (Huang et al., 2023, 2024a) or introduce inaccurate extrapolation errors (Tian et al., 2020), interpolation-based data compressors typically produce quantization codes with a highly concentrated distribution, including very few large values (Liu et al., 2024c). Consequently, applying a short data format to store the quantization codes will accelerate the encoding speed and simplify bit patterns without high overheads for recording outliers.
@@ 5.2.2. Investigation and benchmarking of lossless encoders
In our benchmarking, the evaluated lossless pipelines are collected as follows: First, incorporating Huffman encoding as a ”preprocessor” in the lossless compression pipeline has been a widely adopted scheme in existing compressors (Liang et al., 2023; Liu et al., 2024a, c), so in addition to the standalone lossless modules, we also include a set of Huffman-incorporated variants corresponding to each of them. Second, for the LC framework, we perform preliminary experiments on several datasets and select 8 representative and adaptive pipelines (RRE1, RRE1-RRE2, etc., as shown in Figure 6) with 1/2/3/4 stages (as indicated by Figure 6, pipelines with more stages are not necessary). The benchmarking is conducted on 4 scientific datasets (Zhao et al., 2020a) on a computing platform facilitated with NVIDIA RTX 6000 Ada GPUs.
The compression ratio and overall (compression-decompression) throughput of each lossless pipeline are shown in Figure 6, with Pareto frontiers marked (solutions with throughput less than 25 GiB/s are not included in the frontiers because they bring unacceptable time overheads to the full compression pipeline). Regarding open-source solutions with acceptable compression throughput, we find that the combination of Huffman encoding (hereinafter HF) and the RRE4-TCMS8-RZE1 pipeline achieves high compression ratios on cuSZ-Hi quantization codes. The Zstd (Collet, 2015) compressor in NVIDIA nvCOMP achieves the highest compression ratio. However, not only is this implementation NVIDIA-proprietary, but it also consists of multiple self-contained encoders with coarse component configuration granularity. Consequently, its throughput is significantly low in most scenarios, making it impractical for cuSZ-Hi. The nvCOMP::ANS/Bitcomp and GPULZ may present decent compression ratios in some scenarios, but they show no advantages over the multi-stage LC pipelines. The GDeflate/LZ4/ndzip/Huffman-only lossless solutions exhibit very poor compression ratios and/or throughputs, being infeasible for cuSZ-Hi usage.
According to the benchmarking results and analysis, we adopt the HF-RRE4-TCMS8-RZE1 lossless pipeline in cuSZ-Hi, forming its compression-ratio-preferred mode, cuSZ-Hi-CR. However, Huffman encoding on GPUs also brings a performance bottleneck in the compression framework and lowers the overall compression throughput (Rivera et al., 2022). To this end, regarding Figure 6, cuSZ-Hi integrates an alternative Huffman-encoding-free lossless pipeline, which is the TCMS1-BIT1-RRE1 pipeline with high throughput and decent compression ratio. In §6, we will see that, cuSZ-Hi with this lossless pipeline (the throughput-preferred cuSZ-Hi-TP) also gains very acceptable compression ratios with significantly improved compression throughputs, proving its practical usability.